OpenStudy (anonymous):

[ HOW THIS EQUAL? ] -2sin ^2theta +cos ^2 theta = -3sin^2 theta +1

OpenStudy (anonymous):

\[[HELP] -2\sin ^2\theta +\cos ^2 \theta = -3\sin^2 \theta +1 \]

Parth (parthkohli):

\[\color{#C00}{-2 \sin^2\theta} + \cos^2\theta = \color{#C00}{-3 \sin^2\theta + \sin^2\theta} + \cos^2\theta = -3\sin^2\theta + \color{green}{\sin^2\theta + \cos^2\theta }\]

Parth (parthkohli):

What'd that green part be?

OpenStudy (jhannybean):

So colorful :D

OpenStudy (anonymous):

1

OpenStudy (primeralph):